### v2

@@ Abstract
Accurate rainfall data are crucial for effective climate services, especially in Sub-Saharan Africa, where agriculture heavily depends on rain-fed systems. However, the sparse distribution of rain-gauge networks in the region necessitates reliance on satellite and reanalysis rainfall products (REs) for rainfall estimation. This study evaluated the performance of eight REs — CHIRPS, TAMSAT, CHIRP, ENACTS, ERA5, AgERA5, PERSIANN-CDR, and PERSIANN-CCS-CDR — in Zambia and Ghana using a point-to-pixel validation approach. The analysis encompassed spatial consistency, annual rainfall summaries, seasonal patterns, and rainfall intensity detection across 38 ground stations. Results indicated that no single product performed optimally across all contexts, underscoring the need for application-specific recommendations. All products exhibited a high probability of detection (POD) for dry days in Zambia and northern Ghana (with 70% $<$ POD $<$ 100%, and 60% $<$ POD $<$ 85% respectively), suggesting their potential utility for drought-related studies in these areas. Conversely, all products showed limited skill in detecting heavy and violent rains (with POD close to 0%), rendering them unsuitable for extreme rainfall analysis (such as floods) in the current form. Products integrated with station data (ENACTS, CHIRPS, and TAMSAT) outperformed their counterparts under many contexts, highlighting the importance of calibration with local observations. Bias correction is strongly recommended, as varying levels of biases were evident across different rainfall summaries. A critical area for advancement is extreme rainfall detection. Future research should focus on this aspect.
@@ 1 Introduction
Accurate historical rainfall observations are crucial for the provision of reliable climate services [2, 6, 17, 52], particularly in regions where agriculture and water resources are highly sensitive to climatic variations [43]. In Sub-Saharan Africa, where about 95 percent of agriculture is rain-fed [1, 49], access to precise rainfall data is essential for effective planning and management of agricultural practices. However, weather station and rain-gauge networks in this region are often sparse, both in spatial and temporal coverage, or in some cases, entirely non-existent [41, 2, 11]. This lack of ground-based observations makes it challenging to generate localized climate information, further complicating efforts to mitigate the impacts of climate variability on agricultural productivity and food security.
REs offer a promising solution by providing rainfall estimates with high spatial and temporal resolutions [21, 47]. These products can potentially enhance climate monitoring and forecasting capabilities in regions with limited ground-based observation networks. However, their accuracy and usability must be rigorously validated against available ground-based observations [17, 6, 2] before they can be effectively integrated into climate services. Validation serves as a key step in assessing how closely these estimates align with actual observations, ensuring their reliability for applications such as agriculture, hydrology, and disaster management [21].
@@
In contrast, point-to-pixel validation [45, 2, 47, 17, 55] directly compares REs with point-based rain-gauge observations. One advantage of point-to-pixel validation is its ability to provide a more precise assessment at specific locations, helping to identify discrepancies between satellite estimates and on-the-ground observations. However, a key constraint of this approach is the fact that gauges give point measurements while satellites and reanalysis produce spatial averages, which sometimes struggle with local convective storms and orographic rainfall [47]. Comparing rainfall measured at a specific point to satellite estimates averaged over a large area is inherently problematic [47].
In this study, we performed a comprehensive validation of eight REs in Zambia and Ghana using a point-to-pixel approach. This method was chosen primarily because the available station network was sparse, with no more than one station located within a given satellite or reanalysis grid. The uncertainty of interpolation methods increases significantly as the number of stations within a grid decreases [30, 31, 42]. Additionally, we sought to evaluate the suitability of these REs for providing localized climate information. This is due to the growing demand for gridded data with finer spatial and temporal resolutions, driven by the shift in climate change research from global analyses to more regional and localised investigations [30].
Our validation focused on key aspects such as spatial consistency, annual rainfall summaries, seasonal patterns, and rainfall intensity detection. The importance of validating these REs under diverse conditions has been highlighted in previous studies [17]. Through this multi-dimensional approach, we aimed to provide a robust assessment of the utility of these REs for climate services.
@@ 2.1 Study Area
The validation area covers Zambia and Ghana (Figures 1(a) and 1(b) respectively). Zambia lies between 22 to 34 [33] east of Greenwich, and 8 to 18 south of the equator [35]. The prevailing climate is mainly sub-tropical, witnessing around 95 percent of its total precipitation during the wet season extending from October to April [42] when the Inter-tropical Convergence Zone (ITCZ) is located in the region [28]. Zambia is a landlocked, relatively flat country where rainfall is predominantly the result of convection [42]. The rainfall in this region tends to be more uniform and widespread. The variations in rainfall tend to be tied to the specific agroecological region. In the northern areas, the average annual precipitation exceeds 1200 mm/year, while the southern part experiences less than 700 mm/year annually [35]. The central part receives an annual rainfall of approximately 800 - 1000 mm/year, distributed evenly across the crop growing season [33].
Figure 1: Study area
Ghana is located in West Africa, and lies between latitudes 4.6N and 11N, and longitudes -3.3W and 1.8E [50]. The climate of Ghana is tropical and humid [9]. Two rainy seasons occur in the south from April to July and from September to November (with peaks in June and October respectively), whereas the north has only one rainy season from April to September, with one peak in August [3]. Unlike Zambia, the rainfall pattern in Ghana is quite complex with coastal and orographic influences. The seasonality in the rainfall patterns is brought about by the movement of the Inter-Tropical Convergence Zone (ITCZ) [9, 60]. The country can be broadly categorized under three climatic zones, which are the Savannah zone, the Forest zone, and the Coastal zone [8, 50]. Mean annual rainfall ranges between 900 and 2100 mm, where the south-west has relatively high values while the north, and south-eastern coast have relatively low values [5].
@@
    A unimodal rainfall pattern in the northern part of the continent, with a peak in August, observed in places such as Burkina Faso, Mali, Niger, and northern Ghana.
@@
    A unimodal rainfall pattern in the southern part of the continent, peaking in February, characteristic of regions like Zambia, Mozambique, and Malawi.
@@
    A bimodal rainfall pattern in the mid-belt of the continent, found in areas including southern Ghana and parts of Kenya.
@@ 2.2.1 Station Data
Data from $15$ stations across the country of Zambia, and $23$ stations across Ghana were used in this validation study. The Zambian data was obtained from the Zambia Meteorological Department (ZMD), while the Ghanaian data was obtained from the Ghana Meteorological Agency (GMet). They were daily rain gauge measurements. The data, as were obtained from ZMD and GMet, had longer records going as far back as 1930 for some of the stations. However, 1983 was considered to be the starting period for the stations in order that they coincide with years that most of the REs have available data. Zambian stations with long records (at least 30 years, with the exception of two stations which had 27 and 28 years) were used, and they are scattered unevenly across all the three climatic zones of the country. In the case of Ghana, the 23 stations are synoptic stations [50]. All the data were quality-controlled before use in the study. These quality-control checks included consecutive rain day check, consecutive same rainfall value check, extreme values check, dry months check, and missing values check. This part was mostly done in consultation with the data providers, and took a considerable amount of time as it is a crucial part of the evaluation. Values that did not pass the quality-control checks were replaced with missing data. Table 1 shows the names of the stations, their longitude and latitude information, the period they cover as well as the percentage of complete data within that period.
@@ 2.2.2 Satellite and Reanalysis Data
Eight REs were used in this study, which include six satellite-based and two reanalysis rainfall products. They are the Climate Hazards Group Infrared Precipitation (CHIRP), CHIRP with Stations (CHIRPS) [23, 22], Tropical Application of Meteorology using Satellite data (TAMSAT) [42, 59, 29], the European Centre for Medium-Range Weather Forecasts (ECMWF) v5 (ERA5) [29, 7], fifth-generation reanalysis of ECMWF (AgERA5 hereafter AGERA5) [10, 12], Precipitation Estimation from Remotely Sensed Information using Artificial Neural Networks-Cloud Classification System-Climate Data Record (PERSIANN-CCS-CDR, hereafter PCCSCDR) [54], Precipitation Estimation from Remotely Sensed Information using Artificial Neural Networks - Climate Data Record (PERSIANN-CDR hereafter PCDR) [4], and Enhancing National Climate Services (ENACTS) [16, 18].
@@
Daily rainfall data from REs were extracted at grid points nearest to the station coordinates (Table 1) for comparison with observations. A summary of the REs’ data used are given in Table 2.
@@ 2.4 Comparison of Annual Summaries
Annual summaries (total rainfall, number of rainy days, and mean rain per rainy day) were used to study this aspect of the REs. In the case of Zambia, each year was adjusted to commence in August and end in July in order to capture the whole rain and dry seasons in one cycle.
These summaries were then compared with the observed rainfall summaries. Mean error (ME), percentage bias (PBIAS), the ratio of standard deviations (RSD), and linear correlation coefficient ($r$) were used as continuous evaluation metrics. The ME (1) was used to measure the average difference between REs and gauge observations, the PBIAS (2) was used to assess the systematic bias of REs [41], the r (3) to assess the degree of linear relationship between the REs and gauge observations [62], and the RSD (4) to compare the similarity between the variabilities of the REs and the observed rain rainfall data.
@@ 2.6 Rainfall Intensity
With rainfall intensity classification, we used probability of detection, which is a standard measure for binary classification. It is the likelihood that the REs will correctly detect the same rainfall events on the same days as recorded in the gauge observations. The probability of detection (POD) is given by (8).
@@ 3 Results
The results from the study are presented and discussed under this section.
@@ 3.1 Spatial Consistency of REs
Figure 2: Mean total rainfall (mm/year) of CHIRPS for 1983-2022 in Zambia
@@
Figure 3: Mean total rainfall (mm/year) of PCCSCDR for 1983-2022 in Ghana
@@
Figure 4: Mean total rainfall (mm/year) of the REs for 1983-2022 in Zambia
Figure 5: Mean total rainfall (mm/year) of the REs for 1983-2022 in Ghana
In the case of Ghana, none of the REs displayed spatial inconsistencies and were found to align closely with the country’s climatology. Figure 5c illustrates the mean annual total rainfall for PCDR, which exhibits a pixelated appearance resulting from its coarse resolution of 0.25 by 0.25 (approximately 25 km by 25 km). This pixelation is similar to that observed in ERA5 (and AGERA5 derived from ERA5), as shown in Figure 5a, while CHIRPS and CHIRP (Figures 5b and 5d, respectively) feature a finer resolution of 0.05 by 0.05 (approximately 5 km by 5 km). Despite the pixelation, all REs effectively capture the climatological patterns in Ghana, with the northern regions and the east coast receiving less rainfall, and the central belt and west coast (particularly the Axim area) experiencing significantly higher precipitation. Overall, ENACTS, CHIRPS, CHIRP, and TAMSAT demonstrated a superior spatial consistency with the climatology of Zambia and Ghana.

### v4

@@ Abstract
Accurate rainfall data are crucial for effective climate services, especially in Sub-Saharan Africa, where agriculture depends heavily on rain-fed systems. The sparse distribution of rain-gauge networks necessitates reliance on satellite and reanalysis rainfall products (REs). This study evaluated eight REs—CHIRPS, TAMSAT, CHIRP, ENACTS, ERA5, AgERA5, PERSIANN-CDR, and PERSIANN-CCS-CDR—in Zambia and Ghana using a point-to-pixel validation approach. The analysis covered spatial consistency, annual rainfall summaries, seasonal patterns, and rainfall intensity detection across 38 ground stations. Results showed no single product performed optimally across all contexts, highlighting the need for application-specific recommendations. All products exhibited a high probability of detection (POD) for dry days in Zambia and northern Ghana (70% $<$ POD $<$ 100%, and 60% $<$ POD $<$ 85%, respectively), suggesting their utility for drought-related studies. However, all products showed limited skill in detecting heavy and violent rains (POD close to 0%), making them unsuitable for analyzing such events (e.g., floods) in their current form. Products integrated with station data (ENACTS, CHIRPS, and TAMSAT) outperformed others in many contexts, emphasizing the importance of local observation calibration. Bias correction is strongly recommended due to varying bias levels across rainfall summaries. A critical area for improvement is the detection of heavy and violent rains, with which REs currently struggle. Future research should focus on this aspect.
@@ 1 Introduction
Accurate historical rainfall observations are critical for providing reliable climate services [2, 6, 19, 57], particularly in regions where agriculture and water resources are highly sensitive to climatic variations [48]. In Sub-Saharan Africa, where approximately 95% of agriculture is rain-fed [1, 54], access to precise rainfall data is indispensable for effective agricultural planning and management. For instance, recent floods and droughts have led to poor crop yields or complete crop losses across the continent [37] including Ghana and Zambia, underscoring the need for reliable rainfall information in these places. However, weather station and rain-gauge networks in this region are often sparse, both spatially and temporally, or in some cases, entirely non-existent [46, 2, 13]. This scarcity of ground-based observations poses significant challenges for generating localized climate information, further complicating efforts to mitigate the impacts of climate variability on agricultural productivity and food security.
REs offer a promising solution by providing rainfall estimates with high spatial and temporal resolutions [24, 52]. These REs can potentially enhance climate monitoring and forecasting capabilities in regions with limited ground-based observation networks. However, the REs do not come without potential shortcomings. They are prone to uncertainties arising from indirect estimation methods, reliance on proxy variables, and potential errors in temporal sampling, course resolution, algorithms, or sensor accuracy, which can limit their reliability for climate applications [6, 19, 67]. Therefore their accuracy and usability must be rigorously validated against available ground-based observations [19, 6, 2] before they can be effectively integrated into climate services. Validation serves as a key step in assessing how closely these estimates align with actual observations, ensuring their reliability for applications such as agriculture, hydrology, and disaster management [24].
@@
In contrast, point-to-pixel validation [50, 2, 52, 19, 60, 6] directly compares REs with point-based rain-gauge observations. One advantage of point-to-pixel validation is its ability to provide a more precise assessment at specific locations, helping to identify discrepancies between satellite estimates and on-the-ground observations. However, a key constraint of this approach is the fact that gauges give point measurements while satellites and reanalysis produce spatial averages, which sometimes struggle with local convective storms and orographic rainfall [52]. Comparing rainfall measured at a specific point to satellite estimates averaged over a large area is inherently problematic [52].
While we acknowledge the limitations of the point-to-pixel approach, this study conducted a comprehensive validation of eight REs in Zambia and Ghana using this method. This was chosen primarily because we sought to evaluate the suitability of these REs for providing localized climate information. This is due to the growing demand for gridded data with finer spatial and temporal resolutions, driven by the shift in climate change research from global analyses to more regional and localized investigations [34]. Additionally, the available station network was sparse, with no more than one station located within a given satellite or reanalysis grid. The uncertainty of interpolation methods increases significantly as the number of stations within a grid decreases [34, 35, 47]. The native resolutions of the REs were retained (as also done by [7], and [6]) without remapping them to a common resolution. This decision was intentional, as it allowed us to assess the performance of each RE as they are in their operational forms. While remapping to a common resolution could reduce spatial discrepancies, it might also obscure the inherent characteristics of each RE. Zambia and Ghana were selected because they represent distinct yet complementary rainfall regimes — unimodal in Zambia and northern Ghana, and bimodal in southern Ghana (see Figures 2 and 3) — while also sharing similarities in their reliance on rain-fed agriculture and challenges related to sparse station networks. This combination allowed for a robust evaluation of REs across diverse climatic and geographical contexts.
Our validation focused on key aspects, including spatial consistency, annual rainfall summaries, seasonal patterns, and rainfall intensity detection. The importance of evaluating REs under diverse conditions has been emphasized in previous studies [19]. Through this multi-dimensional approach, we aimed to provide a robust assessment of the utility of these REs for climate services. Besides, this study evaluated a wide range of products, including the Precipitation Estimation from Remotely Sensed Information using Artificial Neural Networks-Cloud Classification System-Climate Data Record (PERSIANN-CCS-CDR; hereafter PCCSCDR) [59] and the Precipitation Estimation from Remotely Sensed Information using Artificial Neural Networks-Climate Data Record (PERSIANN-CDR; hereafter PCDR) [4], which have not been extensively validated across the African continent, especially at our study area. By including these products, we provide new insights to the scientific community. Furthermore, the spatial consistency analysis conducted in this study, an aspect often overlooked in validation efforts, offers an understanding of its significance in assessing the reliability of REs.
@@ 2.1 Study Area
The validation area covers Zambia and Ghana (Figures 1). Zambia lies between 22 to 34 [38] east of Greenwich, and 8 to 18 south of the equator [40]. The prevailing climate is mainly sub-tropical, witnessing around 95 percent of its total precipitation during the wet season extending from October to April [47] when the Inter-tropical Convergence Zone (ITCZ) is located in the region [32]. Zambia is a landlocked, relatively flat country where rainfall is predominantly the result of convection [47]. The rainfall in this region tends to be more uniform and widespread. The variations in rainfall tend to be tied to the specific agroecological region. In the northern areas, the average annual precipitation exceeds 1200 mm/year, while the southern part experiences less than 700 mm/year annually [40]. The central part receives an annual rainfall of approximately 800 - 1000 mm/year, distributed evenly across the crop growing season [38].
Figure 1: Map of study area showing the station locations in Ghana (at the top left) and Zambia (at the top right) and the dominant climatological zones in Africa (at the bottom right) based on Köppen-Geiger climate classification [8] — (1991-2020)
Ghana is located in West Africa, and lies between latitudes 4.6N and 11N, and longitudes -3.3W and 1.8E [55]. The climate of Ghana is tropical and humid [11]. Two rainy seasons occur in the south from April to July and from September to November (with peaks in June and October respectively), whereas the north has only one rainy season from April to September, with one peak around August and September [3]. Unlike Zambia, the rainfall pattern in Ghana is quite complex with coastal and orographic influences. The seasonality in the rainfall patterns is brought about by the movement of the Inter-Tropical Convergence Zone (ITCZ) [11, 66]. The country can be broadly categorized under three climatic zones, which are the Savannah zone, the Forest zone, and the Coastal zone [10, 55]. Mean annual rainfall ranges between 900 and 2100 mm, where the south-west has relatively high values while the north, and south-eastern coast have relatively low values [5].
Figure 2: Ombrothermic diagrams for Tamale (a), Axim (b), chosen to represent the unimodal and bimodal rainfall patterns in the northern and southern parts of Ghana, respectively. The bar charts show the mean monthly total rainfalls while the line plots show the mean maximum temperatures (red) and mean minimum temperatures (black). The Temperature scale $=$ Rainfall scale / 10. Ombrothermic diagrams for other stations are in the Supplementary Materials
Figure 3: Ombrothermic diagram for Livingstone showing the unimodal rainfall pattern in the in Zambia (just as the other stations in the Supplementary Material). The bar charts show the mean monthly total rainfalls while the line plots show the mean maximum temperatures (red) and mean minimum temperature (black). The Temperature scale $=$ Rainfall scale / 10. The year at this station (and all Zambian stations in this work) have been considered to start in August and end in July)
@@
    A unimodal rainfall pattern in the northern part of Sub-Saharan Africa, observed in places such as northern Ghana (as shown on Figure 2a), Burkina Faso, Mali and Niger
@@
    A unimodal rainfall pattern in the southern part of the continent, characteristic of regions like Zambia (Figure 3), Mozambique, and Malawi.
@@
    A bimodal rainfall pattern in the mid-belt of the continent, found in areas including southern Ghana (Figure 2b) and parts of Kenya.
@@ 2.2.1 Station Data
Data from $15$ stations across the country of Zambia, and $23$ stations across Ghana were used in this validation study. The Zambian data was obtained from the Zambia Meteorological Department (ZMD), while the Ghanaian data was obtained from the Ghana Meteorological Agency (GMet). They were daily rain gauge measurements. The data, as were obtained from ZMD and GMet, had longer records going as far back as 1930 for some of the stations. However, 1983 was chosen as the starting point for the stations to align with the years for which most of the REs have available data. Zambian stations with long records (at least 30 years, with the exception of two stations which had 27 and 28 years) were used, and they are scattered unevenly across all the three climatic zones of the country. In the case of Ghana, the 23 stations are synoptic stations [55] and are distributed across the three climatic zones.
A key challenge associated with station data is its potential for errors or inconsistencies. To address this, all station observations underwent rigorous quality-control procedures before being utilized in the study. These procedures included checks for consecutive rain days, repeated identical rainfall values, extreme values, dry months, and missing data. This process was conducted in close collaboration with the data providers and required significant time and effort, given its critical role in ensuring data reliability. Values that failed the quality-control checks were flagged as missing. Only stations with at least 70% non-missing data were included in the analysis. For annual summaries, years with fewer than 355 days of non-missing observations were excluded. Similarly, for seasonal and categorical comparisons, only days with non-missing observations in both the station data and the REs were considered. Table 1 provides information on the stations used, including their names, geographic coordinates, temporal coverage, and the percentage of complete data within the study period.
@@ 2.2.2 Satellite and Reanalysis Data
The eight REs used in this study, which include six satellite-based and two reanalysis rainfall products, were the Climate Hazards Group Infrared Precipitation (CHIRP), CHIRP with Stations (CHIRPS) [26, 25], Tropical Application of Meteorology using Satellite data (TAMSAT) [47, 65, 33], the European Centre for Medium-Range Weather Forecasts (ECMWF) v5 (ERA5) [33, 9], fifth-generation reanalysis of ECMWF (AgERA5 hereafter AGERA5) [12, 14], PCCSCDR [59], PCDR [4], and Enhancing National Climate Services (ENACTS) [18, 20].
@@
We employed a point-to-pixel approach, extracting daily rainfall data from the REs at their grid points closest to the station coordinates for comparison with observational data. However, for certain REs (e.g., PCDR and TAMSAT), data for their nearest grid points to some stations were unavailable due to their location over the sea or largely outside the country’s boundaries. In such cases, the affected stations were excluded from the validation process for those specific REs. Neither the station data were interpolated to grid scales nor were the REs remapped to a common resolution, as the rationale for this decision has been outlined in the introduction. A summary of the REs’ datasets used in this study is provided in Table 2.
@@ 2.4 Comparison of Annual Summaries
Annual summaries (total rainfall, number of rainy days, and mean rain per rainy day) were used to evaluate the performance of the REs. In the case of Zambia, each year was adjusted to commence in August and end in July to capture the entire rain and dry seasons in one cycle.
These summaries were validated against observed rainfall data using a set of statistical metrics. The choice of these metrics was guided by their ability to capture different aspects of RE performance, ensuring a comprehensive evaluation. Specifically, the following metrics were employed:
  • Mean Error (ME) (1): This metric quantifies the average difference between the REs and gauge observations. It provides a straightforward measure of overall bias, indicating whether the REs tend to overestimate (positive values) or underestimate (negative values) rainfall [22]. This is particularly useful for identifying systematic errors in the REs. The ideal value for ME is 0, indicating no bias.
  • Percentage Bias (PBIAS) (2): PBIAS assesses the systematic bias of the REs relative to observed data, expressed as a percentage. It is valuable for understanding the magnitude and direction of bias [46]. PBIAS ranges from $\infty$- ∞ to +$\infty$+ ∞, with 0 indicating no bias, positive values indicating overestimation, and negative values indicating underestimation.
  • Linear Correlation Coefficient ($r$) (3): This metric evaluates the degree of linear relationship between the REs and gauge observations [68]. It ranges from $-1$ to $+1$, where $+1$ indicates a perfect positive linear relationship, $-1$ indicates a perfect negative linear relationship, and 0 indicates no linear relationship. A high correlation indicates that the REs can reliably capture the temporal variability of rainfall.
  • Ratio of Standard Deviations (RSD) (4): RSD compares the variability of the REs with that of the observed rainfall data. An RSD of 1 indicates perfect agreement in variability, values greater than 1 indicate overestimation of variability, and values less than 1 indicate underestimation of variability. This metric is crucial for applications such as flood forecasting and drought monitoring.
These metrics are defined below:
@@ 2.6 Rainfall Intensity
With rainfall intensity classification, we used POD, which is a standard measure for binary classification. It is the likelihood that the REs will correctly detect the same rainfall events on the same days as recorded in the gauge observations. The POD is given by (8).
@@ 3 Results
The results from the study are presented under the following subsections:
@@ 3.1 Spatial Consistency of REs
Figure 4: Mean total rainfall (mm/year) of CHIRPS for 1983-2022 in Zambia for checking the RE’s spatial consistency
@@
Figure 5: Mean total rainfall (mm/year) of PCCSCDR for 1983-2022 in Zambia for checking the RE’s spatial consistency
@@
Figure 6: Mean total rainfall (mm/year) of the REs for 1983-2022 in Zambia for checking the spatial consistency of the REs
Figure 7: Mean total rainfall (mm/year) of the REs for 1983-2022 in Ghana for checking the spatial consistency of the REs
In the case of Ghana, none of the REs displayed spatial inconsistencies and were found to align closely with the country’s climatology. Figure 7c illustrates the mean annual total rainfall for PCDR, which exhibits a pixelated appearance resulting from its coarse resolution of 0.25 by 0.25 (approximately 25 km by 25 km). This pixelation is similar to that observed in ERA5 (and AGERA5 derived from ERA5), as shown in Figure 7a, while CHIRPS and CHIRP (Figures 7b and 7d, respectively) feature a finer resolution of 0.05 by 0.05 (approximately 5 km by 5 km). Despite the pixelation, all REs effectively capture the climatological patterns in Ghana, with the northern regions and the east coast receiving less rainfall, and the central belt and west coast (particularly the Axim area) experiencing significantly higher precipitation. Overall, ENACTS, CHIRPS, CHIRP, and TAMSAT demonstrated a superior spatial consistency with the climatology of Zambia and Ghana.